Groton Construction Products is considering the purchase of one or more of the companies in their supply chain. Calculate the ROI and residual income for each company and indicate if Groton should consider purchasing any of the companies based on the ROI and residual income. Show your calculations for full points.

                                               ____________         Company___________________

                                            _____A_____         _____B_____          ____ C_____  

Sales                                         $3,600,000             $6,000,000                $1,500,000

Net Operating Income                120,000                  220,000                       90,000

Average Operating Assets          1,000,000               2,000,000                       900,000

Minimum required rate of return:       15%                          6%                             10%
Answer #1

ROI = Net operating income/Average operating assets

A = 120,000/1,000,000 = 12%

B = 220,000/2,000,000 = 11%

C = 90,000/900,000 = 10%

Residual income = Net operating income - Required return

A = 120,000 - (1,000,000*15%) = -30,000

B = 220,000 - (2,000,000*6%) = 100,000

C = 90,000 - (900,000*10%) = 0

Comment if you face any issues